OpenStudy (anonymous):

Find the slant height of a cone with surface area 63.30 ft2 and diameter 6.2 ft.

OpenStudy (jack1):

do u know the formula for "total Surface Area of a Cone?"

OpenStudy (jack1):

@dopegyal_ ?

OpenStudy (jack1):

ok, google "surface area of a cone" and get back to me with the formula, and we'll go through it together if u like?

OpenStudy (anonymous):

SA= pi*r*s + pi*r^2 (:

OpenStudy (jack1):

yep now if u think about a cone in terms of Area, it can be split into 2 parts: 

OpenStudy (jack1):

so that's exactly how your equation works SA = pi*r*s + pi*r^2 cone = sector area + circle area

OpenStudy (jack1):

now the question gave you the diameter of the circle, so how do you get the area of a circle if you only know the diameter *remember A (circle) = pi x radius ^2

OpenStudy (jack1):

@dopegyal_ ?

OpenStudy (anonymous):

30.1

OpenStudy (jack1):

cool, so u understand how to calculate that and the relationship between radius and diameter?

OpenStudy (anonymous):

yes (: ty

OpenStudy (jack1):